OpenStudy (anonymous):

The sum of two consecutive odd integers is 36. Find the integers.

OpenStudy (anonymous):

answered this one already. i'll do it again anyway. x+2+x+4=36 2x=30 x=15 15+2=17 15+4=19 answers: 17 and 19.

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

pretty sure this is my second time doing it today lol.

OpenStudy (anonymous):

mine too